2. Find the equation of the tangent line to the curve at the given point. x = 2 - 3 cos , y = 3 + 2 sin a t (-1,3)
Answer #1

x=2-3ws o y=342 Sino 20 a 3500 en = losos Then clore! COS Sino - means Straight live parallel 16 = a it 3= 3 + 2 sino to yuai